2 e у. 6у t 12y - -y +3y е е

is there anyone who can solve this problem?
Answer #1

O chare lan a 3 = 0 = m = 9 + y - byl + 12 yl - 8y = x 0.5 x ₃0 for Homogeneous part y - byl + 12yl - 8y=0 m² smd t2m-8=0 (@ xyl + 3y = 2 > g +3cu - xe for Homogeneous part x8yll + 3x²y=0 & Cauchy Euler} diff. En DCD-1) (0-2) + 3 D (0-1) = 0 D(DWIE 0 0 X X 0 2x3 1 x = -x + x² + x1 (-x) = -2 HgX) ti! 11 0 0 * 2 1(x) - 0 +x (0) = b to 0 4₂ (X) = Х 6 0 0 0 - ger) al llo)